What is the distance between each pair of points:
1.(0,5) and (0,8)
2.(7,0) and (2,0)
3.(4,14) and (4,9)
4.(1,171) and (11,171)
5.(6,-2) and (6,3)

Answer by Boreal(15235)

When the x or y doesn't change the distance is the absolute value (positive number) distance between the other coordinate. If you see that, the problem becomes quite simple.
The distance formula is sqrt{ ((x1-x2)^2+(y1-y2)^2}
For the first, sqrt{(0-0)^2+(5-8)^2}=sqrt (0+(-3)^2)=sqrt(9)=3
The first is just 8-5=3
The second is 7-2=5. If you do 2-7 and get -5, just take the absolute value which makes it +5.
sqrt((7-2)^2+(0-0)^2))=sqrt(25+0)=sqrt(25)=5. Distance is always positive.
The third is 14-9=5
The fourth is abs(1-11)=abs(-10) or 10.
The fifth is 3-(-2) or 5. Watch the sign. You can also look at that as going from -2 to 3, which is a distance of 5.